OpenStudy (anonymous):

hi, can someone help me find the integral (0 to pi/4) of sqrt(1+tan^2 (t)) dt please?

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/4} \sec^2t dt\]

OpenStudy (bahrom7893):

tan(x) from 0 to pie/4

OpenStudy (anonymous):

= 1-0=1

OpenStudy (anonymous):

tan(pi/4)= 1 & tan(0) = 0

OpenStudy (bahrom7893):

lol two people helped u out... fan us both if we helped, thanks!

OpenStudy (anonymous):

I think these guys misread your question. tan(x)^2+1=sec(x)^2 so you have the integral from 0 to pi/4 of [sec(t) dt] The integral of sec(x) is ln(abs[sec(x)+tan(x)])+C where C is a constant and abs[] means absolute value. You can find proof of this integral through a quick google search. You should be able to handle the rest, but I got ln[1+sqrt(2)]